OpenStudy (anonymous):

Which exponent makes the statement true? 1/5^9 = 5^? 9 -9 1/9 -1/9

OpenStudy (anonymous):

\[(\frac{ 1 }{ 5 })^{9} or \frac{ 1 }{ 5^{9} }\]

OpenStudy (anonymous):

\[\frac{1}{5^9 } =5^{-9}\]

OpenStudy (anonymous):

So, -9 is the correct Answer

OpenStudy (anonymous):

Don't just give answers Dpasingh.

OpenStudy (anonymous):

@malical Do you think that it is so. Pls check the page.

OpenStudy (anonymous):

@KaterinaGilbert

OpenStudy (anonymous):

What is the page?

OpenStudy (anonymous):

@malical where you are posting